OpenStudy (ksaimouli):

solve for x cos^2x-sin^2x=sin x

OpenStudy (jackellyn):

Use the identity (sinx)^2+(cosx)^2=1 or (cosx)^2=1-(sinx)^2 So: 1-(sinx)^2-(sinx)^2=sin(x) 1-2(sinx)^2=sin(x) 0=2(sinx)^2+sin(x)-1 sinx=[-1+sqrt(1+8)]/4 or [-1-sqrt(1+8)]/4 sinx=[-1+3]/4 or [-1-3]/4 sinx=2/4 or -4/4 sinx=1/2 or -1 If sinx=0.5 x=pi/6+2piK or x=5pi/6+2piK If sinx=-1 x=-pi/2+2piK All the possible solutions are actually x=pi/6+(2/3)piK And -pi<x<=pi The solutions are: x= -pi/2, pi/6, 5pi/6